Reduce each equation to standard form. Then find the coordinates of the center, the foci, the ends of the major and minor axes, and the ends of each latus rectum. Sketch the curve

(a) x2 + 4y2 + 6x + 16y + 21 = 0

(b) 16x2 + 4y2 + 32x − 16y − 32 = 0

(c) 4x2 + 8y2 − 4x − 24y − 13 = 0


1
Expert's answer
2021-08-19T13:11:28-0400

(a)


"x^2 + 4y^2 + 6x + 16y + 21 = 0"

"(x+3)^2+4(y+2)^2=4"

"\dfrac{(x+3)^2}{4}+\dfrac{(y+2)^2}{1}=1"

Center: "(-3, -2)"

"c=\sqrt{a^2-b^2}=\sqrt{4-1}=\sqrt{3}"

Foci: "(-3-\sqrt{3}, -2), (-3+\sqrt{3}, 2)"

Vertices: "(-5, -2), (-1, -2)"

Co-vertices: "(-3, -3), (-3, -1)"

First latus rectum: "x=-3-\sqrt{3}"

Second latus rectum: "x=-3+\sqrt{3}"

The length of the latera recta: "1"

"(-3-\sqrt{3}, -2.5), (-3-\sqrt{3}, -1.5)"

"(-3+\sqrt{3}, -2.5), (-3+\sqrt{3}, -1.5)"




(b)


"16x^2 + 4y^2 + 32x \u2212 16y \u2212 32 = 0"

"16x^2+32x+16+4y^2-16y+16=64"

"16(x+1)^2+4(y-2)^2=64"

"\dfrac{(x+1)^2}{4}+\dfrac{(y-2)^2}{16}=1"

Center: "(-1, 2)"

"c=\sqrt{a^2-b^2}=\sqrt{16-4}=2\sqrt{3}"

Foci: "(-1,2-2\sqrt{3}), (-1,2+2\sqrt{3})"

Vertices: "(-1, -2), (-1, 6)"

Co-vertices: "(-3,2), (1, 2)"

First latus rectum: "y=2-2\sqrt{3}"

Second latus rectum: "y=2+2\sqrt{3}"

The length of the latera recta: "1"

"(0, 2-2\sqrt{3}), (0,2+2\sqrt{3})"

"(1,2-2\sqrt{3}), (1, 2+2\sqrt{3})"




(c)


"4x^2 + 8y^2 \u2212 4x \u2212 24y \u2212 13 = 0"


"4x^2-4x+1+8y^2-24y+18=32"

"4(x-\dfrac{1}{2})^2+8(y-\dfrac{3}{2})^2=32"

"\dfrac{(x-\dfrac{1}{2})^2}{8}+\dfrac{(y-\dfrac{3}{2})^2}{4}=1"

Center: "(\dfrac{1}{2},\dfrac{3}{2})"


"c=\sqrt{a^2-b^2}=\sqrt{8-4}=2"


Foci: "(-\dfrac{3}{2},\dfrac{3}{2}), (\dfrac{5}{2},\dfrac{3}{2})"


Vertices: "(\dfrac{1}{2}-2\sqrt{2},\dfrac{3}{2}),(\dfrac{1}{2}+2\sqrt{2},\dfrac{3}{2})"


Co-vertices: "(\dfrac{1}{2},-\dfrac{1}{2}), (\dfrac{1}{2},\dfrac{7}{2})"


First latus rectum: "x=-\dfrac{3}{2}"

Second latus rectum: "x=\dfrac{5}{2}"

The length of the latera recta: "2\sqrt{2}"

"(-\dfrac{3}{2}, \dfrac{3}{2}-2\sqrt{2}), (-\dfrac{3}{2},\dfrac{3}{2}+2\sqrt{2})"


"(\dfrac{5}{2},\dfrac{3}{2}-2\sqrt{2}), (\dfrac{5}{2}, \dfrac{3}{2}+2\sqrt{2})"